# Explaining the Implicit Neural Canvas: Connecting Pixels to Neurons by Tracing their Contributions

### v1

@@ Abstract
The many variations of Implicit Neural Representations (INRs), where a neural network is trained as a continuous representation of a signal, have tremendous practical utility for downstream tasks including novel view synthesis, video compression, and image superresolution. Unfortunately, the inner workings of these networks are seriously under-studied. Our work, eXplaining the Implicit Neural Canvas (XINC), is a unified framework for explaining properties of INRs by examining the strength of each neuron’s contribution to each output pixel. We call the aggregate of these contribution maps the Implicit Neural Canvas and we use this concept to demonstrate that the INRs which we study learn to “see” the frames they represent in surprising ways. For example, INRs tend to have highly distributed representations. While lacking high-level object semantics, they have a significant bias for color and edges, and are almost entirely space-agnostic. We arrive at our conclusions by examining how objects are represented across time in video INRs, using clustering to visualize similar neurons across layers and architectures, and show that this is dominated by motion. These insights demonstrate the general usefulness of our analysis framework. Our project page is available here.
@@ 1 Introduction
In spite of their immense utility, the inner workings of INRs remain relatively under-explored. While some studies shed light on MLP-based INRs [51], NeRF [52], and the impact of hypernetworks [10, 51], the hidden details of non-coordinate methods such as NeRV [7] are even more opaque. This is not completely unique to INRs – explainability and interpretability are still massive challenges for other areas in vision as well [17, 2, 1, 34]. Nevertheless, there is a significant gap for understanding INRs, considering that popular analysis techniques such as NetDissect [3] and GradCAM [42] do not transfer to INRs in straightforward ways. Even analysis work for generative networks tends to focus on control and manipulation [21], and the coordinate-to-signal paradigm of INR is distinct enough from the generative noise-to-signal paradigm that most analysis work cannot be trivially extended [33].
@@
Motion Drives Contribution Change Over Time We show how NeRV’s handle categories of objects and objects in motion, demonstrating that contributions to a given object remain relatively constant over time. This is true even for objects in motion, and fluctuations in our neural canvas across time for videos prove that motion drives changes in which pixels each neuron contributes to.
@@ 2 Related Work
Implicit Neural Representation is a way to represent and compactly encode a variety of high resolution signals. The key idea is to map a set of coordinates to a specific signal, by employing a neural network as the approximating function [44, 31, 7, 49, 41]. SIREN [44] utilizes periodic activation functions in MLPs to fit and map complicated signals, including images, 3D shapes and videos using very small networks. COIN [13], the first image-specific INR method, uses implicit neural representation for image compression tasks and COIN++ [14] extends this work to encode multiple images. In the challenging realm of videos, NeRV [7] is the first method to scale video compression using image-wise implicit representation. It uses convolution layers in addition to MLPs and outputs all RGB values of a frame given the positional embedding of frame index $t$ as input. HNeRV [9] treats the frame itself as the index embedding by introducing a small encoder, thus improving performance over index-based methods. NIRVANA [29] harnesses spatio-temporal redundancies in videos by treating them as groups of frames and fitting distinct networks to each group, enabling it to encode videos of varying lengths and resolutions. Meta-learning approaches try to address long per-item fitting times by learning a hypernetwork for predicting network weights [10, 23]. [28] further expand on the utility of INR by proposing a method for encoding weights as vectors for a wide variety of downstream tasks.
@@
Figure 2: (left) We dissect MLP-based INRs by aggregating their activations (weights multiplied by previous layer outputs) for each pixel at each neuron. (right) We extend this core idea of pixel-to-neuron mapping for the CNN-based INR, NeRV, by computing intermediate feature maps that are not yet summed on the input dimension. For layers prior to the head, we also account for the PixelShuffle and apply a filter to account for the subsequent layer’s kernels, which propagate that neuron’s contribution to neighboring pixels. We provide more details in the appendix.
@@
Interpreting and Visualizing Deep Neural Networks is a longstanding field of study in machine learning. A large number of techniques aim to understand the internal representations of convolutional neural networks and deep networks in general. [20] analyzes effects of various object characteristics on detection performance. Grad-CAM [42] aims to provide visual explanations for the decisions of CNNs. Using the gradients flowing into the final convolutional layer from a target concept, Grad-CAM and related subsequent works [12, 6, 15] generate a coarse localization map that highlights crucial image regions contributing to the prediction of the concept, and [22] proposes a method for automatically captioning these regions to describe image features. Network Dissection [3] considers each unit in a network as a concept detector and compares their activity with human-interpretable pattern-matching tasks such as the detection of object classes. GAN Dissection [4] extends these ideas to visualize and understand GANs at the unit, object, and scene-level. Other works characterize neuron behavior with natural language [19], by combining NetDissect-style concepts [32, 48, 25], or in terms of how they behave in conjunction with other closely-related neurons [37]. [26] creates a comprehensive test methodology that systematically perturbs model inputs and monitors the effect on model predictions, helping identify potential instabilities. These works are insightful for their target domains, but do not directly apply to INR. So, our work proposes to analyze the relationships between various parameters in MLP- and CNN-based INR networks and the spatio-temporal locations of the reconstructed output.
@@ 3.1 Dissecting Multi-Layer Perceptrons
In this study, as a representative method for MLP-based INRs, we use a simple Fourier Feature Network (FFN) [47] with a Random Fourier Matrix [39] as the positional encoding, followed by linear layers. We represent the MLP as $\theta$ and the $n$ linear layers as $\{l_{1},l_{2},\dots,l_{n}\}$. At each MLP layer, the input pixels are operated upon independently, without any spatial rearrangement or sharing of information between the pixels. Thus, we can directly obtain a mapping from the neurons in a layer to the $h\times w$ spatial locations (pixels) in the output image.
For each neuron, $\eta$, we compute a mapping of how its output contributes to all neurons in the next layer (or the output channels, for the last layer). Consider a layer, $l$, with a weight matrix $W$ of shape $m\times n$. We use $i$ and $j$ to denote the weight connecting $i$-th neuron in the current layer to $j$-th neuron in the next layer, which are also the indices of the current neuron of interest, $\eta_{i}$, and some neuron in the next layer, $\eta_{j}$, respectively. When we map a neuron, $\eta_{i}$, to the MLP INR’s output, we are thus forming a mapping from the weights $\{w_{i1},w_{i2},\dots,w_{in}\}$ to the output. Now consider the image the INR represents, $I$, of shape $h\times w\times 3$. Let us consider that the INR takes pixel location ${x_{k},y_{k}}$ as input, and neuron $\eta_{i}$ produces an output scalar, $o$. We compute its contribution to pixel ${x_{k},y_{k}}$ as $\sum_{j=0}^{n}o\cdot w_{ij}$. When we compute this for all pixels, we obtain a map of shape $h\times w$ for $\eta_{i}$, which we refer to as the contribution map for $\eta_{i}$.
@@ 3.2.1 Head Layer
We begin by analyzing the contribution from kernels in the convolutional head layer to the output RGB image pixels. Consider the head layer, $l_{h}$, with input $v_{\text{in}}$ of shape $h\times w\times ch_{\text{in}}$. We thus have a set of $ch_{\text{in}}*ch_{\text{out}}$ convolutional kernels, meaning $ch_{\text{in}}$ kernels per each R, G, B channel of the output, and thus, $ch_{\text{in}}*3$ kernels that can potentially contribute to each output pixel. We obtain the individual contribution of each kernel by passing each kernel in $l_{h}$ over $v_{\text{in}}$ and storing the outputs separately, yielding a set of features maps, $h\times w\times(ch_{\text{in}}*ch_{\text{out}})$. To get the true output contribution, and thus the final contribution maps for all $ch_{\text{in}}*3$ neurons, we pass these feature maps through the $l_{h}$ activation (e.g. tanh).
3.2.2 PixelShuffle
@@
All NeRV layers except the head layer perform some upsampling using PixelShuffle [43]. PixelShuffle upsamples feature maps by moving elements from the channel dimension into the spatial dimension. Let the number of input channels of the convolution preceding the PixelShuffle in layer $i$ be $ch_{i,\text{in}}$ and the number of output channels be $ch_{i,\text{out}}$. For upsampling factor $r$, the number of output channels from the PixelShuffle, which is the same as the number of input channels to the $(i+1)^{th}$ layer, is computed as
$ch_{i+1,\text{in}}=\frac{ch_{i,\text{out}}}{r^{2}}$ (1)
@@
We require a contribution map of shape $h\times w\times(ch_{i,\text{in}}*ch_{i,\text{out}})$. To obtain this, we first follow the same approach to the head layer and obtain a mapping of shape $\frac{h}{r}\times\frac{w}{r}\times ch_{i,\text{in}}\times ch_{i,\text{out}}$. Considering the penultimate NeRV layer for the sake of simplicity, we can apply PixelShuffle to this downsampled contribution map and obtain the map of size $h\times w\times ch_{i,\text{in}}\times ch_{(i+1),\text{in}}$. However, the contributions of distinct output kernels $ch_{i,\text{out}}$ are now no longer separated since they are moved into the spatial dimension.
@@
To resolve this, we repeat the operation $r^{2}$ times, yielding a new map of size $r^{2}\times h\times w\times ch_{i,\text{in}}\times ch_{(i+1),\text{in}}$. This is equivalent to a mapping of $h\times w\times ch_{i,\text{in}}\times ch_{i,\text{out}}$ by moving the $r^{2}$ dimension to the end and substituting according to Equation 1. While this map has the dimensions we require for further processing by downstream layers, we must take care that each non-overlapping $r\times r$ block contains samples from the same filter in order to preserve each filter’s contribution. If nearby samples were from different filters, as in the traditional PixelShuffle, then the contribution of different filters would be mixed and aggregated in future layers. We can conveniently accomplish this with a simple re-arranging of the repeated channels followed by masking such that each $r\times r$ block contains only one non-zero element (to prevent overcounting the contribution of each kernel). The final result is then passed through the activation layer before proximity correction.
@@ 3.2.3 Proximity Correction
Layers before the head layer do not contribute directly to pixels. Instead, their outputs are the inputs to some later set of convolution kernels. While we account for the upsampling with how we handle PixelShuffle, we still must address the fact that these kernels act on $k\times k$ neighborhoods at each filter map. So, we use an aggregation filter to account for the fact that a given point in some intermediate feature map contributes to a $k\times k$ neighborhood of points in the next layer’s feature maps. In the case of the penultimate layer, this is a simple $k\times k$ box filter. For the sake of clarity, we reserve further details for our code.
@@ 4.1 Dataset and Training Details
Figure 3: The implicit neural canvas. We show representative example contribution maps for various layers of FFN [47] and NeRV [7]. Notice how early FFN layers manifest strong fourier patterns, and how the last layers of both FFN and NeRV tend to resemble the image, with NeRV layers being reminiscent of classical image processing filters.
@@
To analyze the relationship between neurons and pixels in encoded signals, we use videos from the Cityscapes-VPS [24] and VIPSeg [30] video panoptic segmentation datasets. To select a few videos from these datasets, we look for examples containing objects of diverse shapes, sizes and categories, to afford diversity in analysis. Table 1 provides the number of instances per frame in each dataset, the percentage of annotated frames and other information. Having instance and object category level information enables us to answer questions pertaining to whether INRs understand semantics and instances and also helps analyze the distribution of representations for objects of different types. Further, it would help us determine whether the groups of parameters attending to the same objects remains consistent across frames, whether the model omits representing small objects, and investigate other important properties. We train our FFNs and NeRVs on downsampled images, to keep the model and contribution map sizes small. Specifically, we resize and crop all videos to $128\times 256$ resolution. We choose model sizes and training times such that bits-per-pixel and reconstruction quality are roughly equivalent across models and video frames. We provide more details in the appendix.
@@ 4.2 Contributions Correlate with Colors and Edges
Figure 3 shows some example contribution maps for neurons at different layers for both MLP-based and CNN-based INRs. FFN and NeRV head layer neurons seem to learn some form of whole scene representation, and the CNN penultimate layer has similar characteristics. The contribution maps of the FFN layers suggest a progression from representing Fourier features in early layers, to representing the image in question, with Fourier artifacts, in the last layer. For both the NeRV layers, there seem to be a variety of features captured by various neurons, including edges, textures, colors, and depth. These maps show that the behavior of a neuron is not simply to represent a pixel or set of pixels, instead, it learns some attributes of the scene.
@@
Figure 6: Pixels per neuron. We compute the pixels activated for each neuron, and plot them for each layer, with each layer sorted independently by order of increasing pixels, to show the trends in how many pixels each neuron is responsible for, for frames from two different videos.
We explore these behaviors quantitatively by aggregating neuron contributions to pixels, first in terms of instances, and then also by clustering pixels by color value (RGB Clusters), clustering on Gabor filter features (Gabor Clusters, details in appendix), and clustering pixels spatially (taking equal size rectangular gridcells). We hypothesize that for a type of pixel clustering to be meaningful for neurons, each neuron should tend to have high activation in some clusters, and little to no activation in others. Quantitatively, the variance of contributions across clusters would be high. However, using raw variance of contributions would give massive bias to clusterings that form large and small regions, where some contributions would be trivially small. So instead, we first compute the expected contribution for each cluster, which is the contribution over the whole image, normalized by size of the given cluster.
@@
We plot the result of performing this computation for Instances, RGB, Gabor, and gridcell clusters, across all neurons, for all three layers of the FFN and the last two layers of NeRV, in Figure 4. This computation reveals quantitatively what Figure 3 hints at – there is relatively little meaningful correlation between contributions and space, as seen in the low variances of gridcell representations. Instead there is in general a surprisingly high variance for neurons when we group their contributions using instance masks, which suggests that they learn some type of object semantics. Given that the variances for RGB and Gabor clusters are also quite high, we suggest that these are low-level semantics, a combination of color and edge features. There is also an interesting trend with the NeRV Block 3, where it seems to have some meaningful correlation with gridcells. However, we note that the top 25% of variances for Instances, RGB, and Gabor, are almost all higher than even the highest gridcell variance, and therefore we believe the intuition holds.

### v2

@@ Abstract
The many variations of Implicit Neural Representations (INRs), where a neural network is trained as a continuous representation of a signal, have tremendous practical utility for downstream tasks including novel view synthesis, video compression, and image super-resolution. Unfortunately, the inner workings of these networks are seriously under-studied. Our work, eXplaining the Implicit Neural Canvas (XINC), is a unified framework for explaining properties of INRs by examining the strength of each neuron’s contribution to each output pixel. We call the aggregate of these contribution maps the Implicit Neural Canvas and we use this concept to demonstrate that the INRs we study learn to “see” the frames they represent in surprising ways. For example, INRs tend to have highly distributed representations. While lacking high-level object semantics, they have a significant bias for color and edges, and are almost entirely space-agnostic. We arrive at our conclusions by examining how objects are represented across time in video INRs, using clustering to visualize similar neurons across layers and architectures, and show that this is dominated by motion. These insights demonstrate the general usefulness of our analysis framework.
**footnotetext: Equal contribution
@@ 1 Introduction
In spite of their immense utility, the inner workings of INRs remain relatively under-explored. While some studies shed light on MLP-based INRs [55], NeRF [56], and the impact of hypernetworks [10, 55], the hidden details of non-coordinate methods such as NeRV [7] are quite opaque. This is not completely unique to INRs – explainability and interpretability are still massive challenges for other areas in vision as well [18, 2, 1, 38]. Nevertheless, there is a significant gap for understanding INRs, considering that popular analysis techniques such as NetDissect [3] and GradCAM [46] do not transfer to INRs in straightforward ways. Even analysis work for generative networks tends to focus on control and manipulation [23], and the coordinate-to-signal paradigm of INR is distinct enough from the generative noise-to-signal paradigm that most analysis work cannot be trivially extended [37].
@@
Motion Drives Contribution Change Over Time We show how NeRVs handle objects and motion, demonstrating that contributions to a given object remain fairly constant over time. This is true even with motion, and fluctuations in our neural canvas across time for videos prove that motion drives changes in neuron contributions for NeRV.
@@ 2 Related Work
Implicit Neural Representation is a way to represent and compactly encode a variety of high resolution signals. The key idea is to map a set of coordinates for a specific signal to a function, by employing a neural network as the continuous approximating function [48, 35, 7, 53, 45]. SIREN [48] utilizes periodic activation functions in MLPs to fit and map complicated signals, including images, 3D shapes and videos using very small networks. COIN [13], the first image-specific INR method, uses implicit neural representation for image compression tasks and COIN++ [14] extends this work to encode multiple images. In the challenging realm of videos, NeRV [7] is the first method to scale video compression using image-wise implicit representation. It uses convolution layers in addition to MLPs and outputs all RGB values of a frame given the positional embedding of frame index $t$ as input. Other methods rely on MLPs [27, 32]. Others [8, 9, 19, 58] compute embeddings on the actual frames by introducing a small encoder, thus improving performance over index-based methods. Meta-learning approaches try to address long per-item fitting times by learning a hypernetwork for predicting network weights [10, 25]. Others propose encoding weights as vectors for a wide variety of downstream tasks [31].
@@
Figure 2: (left) We dissect MLP-based INRs by aggregating their activations (weights multiplied by previous layer outputs) for each pixel at each neuron. (right) We extend this core idea of pixel-to-neuron mapping for the CNN-based INR, NeRV, by computing intermediate feature maps that are not yet summed on the input dimension. For layers prior to the head, we also account for the PixelShuffle and apply an aggregation filter to account for the subsequent layer’s kernels, which propagate that neuron’s contribution to neighboring pixels. To compute contributions for a given layer, we simply perform the shown steps in sequence for that layer and each subsequent layer.
@@
Interpreting and Visualizing Deep Neural Networks is a longstanding field of study in machine learning. A large number of techniques aim to understand the internal representations of convolutional neural networks and deep networks in general. Hoiem et al. analyzes effects of various object characteristics on detection performance. Grad-CAM [46] aims to provide visual explanations for the decisions of CNNs. Using the gradients flowing into the final convolutional layer from a target concept, Grad-CAM and related subsequent works [12, 6, 15] generate a coarse localization map that highlights crucial image regions contributing to the prediction of the concept, and [24] proposes a method for automatically captioning these regions to describe image features. Network Dissection [3] considers each unit in a network as a concept detector and compares their activity with human-interpretable pattern-matching tasks such as the detection of object classes. GAN Dissection [4] extends these ideas to visualize and understand GANs at the unit, object, and scene-level. Other works characterize neuron behavior with natural language [21], by combining NetDissect-style concepts [36, 52, 28], or in terms of how they behave in conjunction with other closely-related neurons [41]. Lindsay and Bau creates a comprehensive test methodology that systematically perturbs model inputs and monitors the effect on model predictions, helping identify potential instabilities. These works are insightful for their target domains, but do not directly apply to INR. So, we propose to analyze the relationships between various parameters in MLP- and CNN-based INR networks and the spatio-temporal locations of the reconstructed output.
@@ 3.1 Dissecting Multi-Layer Perceptrons
In this work, as a representative method for MLP-based INRs, we use a Fourier Feature Network (FFN) [51] with a Random Fourier Matrix [43] as the positional encoding, followed by linear layers. At each MLP layer, the input pixels are operated upon independently, without any spatial rearrangement or sharing of information between the pixels. Thus, we can obtain a mapping from the neurons in a layer to the $h{\times}w$ spatial locations (pixels) in the output image.
For each neuron, $\eta$, we compute a mapping of how, given its input, it contributes to every pixel in the final image. Consider a layer, $l$, with a weight matrix $W$ of shape $m{\times}n$. We use $i$ and $j$ to denote the weight connecting $i$-th neuron in the previous layer to $j$-th neuron in the current layer, which are also the indices of the current neuron of interest, $\eta_{j}$, and some neuron in the previous layer, $\eta_{i}$, respectively. Now consider the image the INR represents, $I$, of shape $h{\times}w{\times}3$. Let us consider that the INR takes pixel location ${x_{k},y_{k}}$ as input, and neuron $\eta_{j}$ receives an input vector v. We compute its contribution to pixel ${x_{k},y_{k}}$ as $\sum_{i=1}^{m}\textbf{v}_{i}\cdot w_{ij}$. When we compute this for all pixels, we obtain a map of shape $h{\times}w$ for $\eta_{j}$, which we refer to as the contribution map for $\eta_{j}$.
@@ 3.2.1 Head Layer
We begin by analyzing the contribution from kernels in the convolutional head layer to the output RGB image pixels. Consider the head layer, $l_{h}$, with input $v_{\text{in}}$ of shape $h{\times}w{\times}\text{ch}_{\text{in}}$ ($\text{ch}_{\text{in}}$ and $\text{ch}_{\text{out}}$ are the number of input and output channels, respectively). We thus have a set of $\text{ch}_{\text{in}}*\text{ch}_{\text{out}}$ convolutional kernels, meaning $\text{ch}_{\text{in}}$ kernels per each R, G, B channel of the output, and thus, $\text{ch}_{\text{in}}*3$ kernels that can potentially contribute to each output pixel. We obtain the individual contribution of each kernel by passing each kernel in $l_{h}$ over $v_{\text{in}}$ and storing the outputs separately, yielding a set of feature maps, $h{\times}w{\times}(\text{ch}_{\text{in}}*\text{ch}_{\text{out}})$. To get the true output contribution, and thus the final contribution maps for all $\text{ch}_{\text{in}}*3$ neurons, we pass these feature maps through the $l_{h}$ activation (e.g. tanh).
3.2.2 Prior NeRV Blocks
@@
All NeRV layers except the head layer perform some upsampling using PixelShuffle [47]. PixelShuffle upsamples feature maps by moving elements from the channel dimension into the spatial dimension. Let the number of input channels of the convolution preceding the PixelShuffle in layer $i$ be $\text{ch}_{i,\text{in}}$ and the number of output channels be $\text{ch}_{i,\text{out}}$. For upsampling factor $r$, the number of output channels from the PixelShuffle, which is the same as the number of input channels to the $(i+1)^{\text{th}}$ layer, is computed as
$\text{ch}_{i+1,\text{in}}=\frac{\text{ch}_{i,\text{out}}}{r^{2}}$ (1)
@@
We require a contribution map of shape $h{\times}w{\times}(\text{ch}_{i,\text{in}}*\text{ch}_{i,\text{out}})$. To obtain this, we first follow the same approach to the head layer and obtain a mapping of shape $\frac{h}{r}{\times}\frac{w}{r}{\times}\text{ch}_{i,\text{in}}{\times}\text{ch}_{i,\text{out}}$. Considering the penultimate NeRV layer for the sake of simplicity, we can apply PixelShuffle to this downsampled contribution map and obtain the map of size $h{\times}w{\times}\text{ch}_{i,\text{in}}{\times}\text{ch}_{(i+1),\text{in}}$. However, the contributions of distinct output kernels $\text{ch}_{i,\text{out}}$ are now no longer separated since they are moved into the spatial dimension.
@@
To resolve this, we repeat the operation $r^{2}$ times, yielding a new map of size $h{\times}w{\times}\text{ch}_{i,\text{in}}{\times}\text{ch}_{(i+1),\text{in}}{\times}r^{2}$, which has the same dimensions as a mapping of shape $h{\times}w{\times}\text{ch}_{i,\text{in}}{\times}\text{ch}_{i,\text{out}}$, if we substitute according to Equation 1. While this map has the dimensions we require for further processing by downstream layers, we must take care that each non-overlapping $r{\times}r$ block contains samples from the same filter in order to preserve each filter’s contribution. If nearby samples were from different filters, as in the traditional PixelShuffle, then the contribution of different filters would be mixed and aggregated in future layers. We can conveniently accomplish this with a simple re-arranging of the repeated channels followed by masking such that each $r{\times}r$ block contains only one non-zero element (to prevent overcounting of the contribution of each kernel). The final result is then passed through the activation layer before proximity correction.
Layers before the penultimate layer differ only slightly. For the layer in question, we treat it exactly like the penultimate layer. However, the result is not at the target output resolution. So, we must perform the subsequent layers’ corresponding operations, with one minor adjustment – we use nearest neighbor upsampling, instead of PixelShuffle, for all layers after the initial layer. This is because the spatial correspondences of the PixelShuffle are specific to the feature map structure of the given subsequent layer, and have no meaning for the earlier layer feature maps.
@@ 3.2.3 Proximity Correction
Layers before the head layer do not contribute directly to pixels. Instead, their outputs are the inputs to some later set of convolution kernels. While we account for the upsampling with how we handle PixelShuffle, we still must address the fact that the subsequent kernels act on $k{\times}k$ neighborhoods at each filter map. So, we use an aggregation filter to account for the fact that a given point in some intermediate feature map contributes to a $k{\times}k$ neighborhood of points in the next layer’s feature maps. This is a simple $k{\times}k$ box filter for each layer.
@@ 4.1 Dataset and Training Details
Figure 3: The implicit neural canvas. We show representative example contribution maps for various layers of FFN [51] and NeRV [7]. Notice how early layer FFN neurons manifest strong Fourier patterns, and how the last layers of NeRV tend to resemble the image, with NeRV head layer neurons being reminiscent of classical image processing filters.
@@
To analyze the relationship between neurons and pixels in encoded signals, we use videos from the Cityscapes-VPS [26] (train and val) and VIPSeg [34] video panoptic segmentation datasets. Table 1 provides the number of instances per frame in each dataset, the percentage of annotated frames and other information. To select a few videos from these datasets, we look for examples containing objects of diverse shapes, sizes and categories, to afford diversity in analysis. Having instance and object category level information enables us to answer questions pertaining to whether INRs understand semantics and instances and also helps analyze the distribution of representations for objects of different types. Further, it would help us determine whether the groups of parameters attending to the same objects remains consistent across frames, whether the model omits representing small objects, and investigate other important properties. We train our FFNs and NeRVs on downsampled images, to keep the model and contribution map sizes small. Specifically, we resize and crop all videos to $128{\times}256$ resolution. We choose model sizes and training times such that bits-per-pixel and reconstruction quality are roughly equivalent across models and video frames. We provide more details in the appendix.
@@ 4.2 Contributions Correlate with Colors and Edges
Figure 3 shows some example contribution maps for neurons at different layers for both MLP-based and CNN-based INRs. FFN and NeRV head layer neurons seem to learn some form of whole scene representation, and the CNN penultimate layer has similar characteristics. The contribution maps of the FFN layers suggest a progression from representing Fourier features in early layers, to representing more of the image in question, with Fourier artifacts, in the last layer. For the NeRV layers, there seem to be a variety of features captured by various neurons, including edges, textures, colors, and depth. The behavior of a neuron, in both networks, is not simply to represent some pixels, instead, it learns some low-level scene attributes.
@@
Figure 6: Pixels per neuron. We compute the pixels activated for each neuron in each layer, with each layer sorted independently by order of increasing number of pixels, for frames from two videos.
We explore these behaviors quantitatively by aggregating neuron contributions to pixels, first in terms of instances, and then also by clustering pixels by color value (RGB Clusters), clustering on Gabor filter features [16, 33] (Gabor Clusters, details in appendix), and clustering pixels spatially (taking equal size rectangular gridcells). We hypothesize that for a type of pixel clustering to be meaningful for neurons, each neuron should tend to have high activation in some clusters, and little to no activation in others. Quantitatively, the variance of contributions across clusters would be high. However, using raw variance of contributions would give massive bias to clusterings that form large and small regions, where some contributions would be trivially small. So instead, we first compute the expected contribution for each cluster, which is the contribution over the whole image, normalized by size of the given cluster.
@@
We plot the result of performing this computation for instances, RGB, Gabor, and gridcell clusters, across all neurons, in Figure 4. This computation reveals quantitatively what Figure 3 hints at – there is relatively little meaningful correlation between contributions and space, as seen in the low variances of gridcell representations. Instead there is in general a surprisingly high variance for neurons when we group their contributions using instance masks, which suggests that they learn some type of object semantics. Given that the variances for RGB and Gabor clusters are also quite high, we suggest that these are low-level semantics, a combination of color and edge features. There is also an interesting trend with the NeRV Block 3, where it seems to have some meaningful correlation with gridcells. However, we note that the top 25% of variances for Instances, RGB, and Gabor, are almost all higher than even the highest gridcell variance, and therefore we believe the intuition holds.